OpenStudy (anonymous):

Suppose that $1300 is initially invested in an account at a fixed interest rate, compounded continuously. Suppose also that, after two and a half years, the amount of money in the account is $1367. Find the interest rate per year

razor99 (razor99):

hey u there

razor99 (razor99):

hello?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i figured out the answer already its 26.80% :)

razor99 (razor99):

wait how did u get that

OpenStudy (anonymous):

1300-1367=67 67/30=2.23333 (ans)*12=26.80 check... 26.8*2.5=67 1300+67 from 2.5 years =1367

OpenStudy (barrycarter):

Umm, it's actually about 2.013%: http://fwd4.me/0udS Remember, this is continuous compounding. Also, 26.8% of 1300 is 348.40, which is way more than 67.